# High-dimensional structure underlying individual differences in naturalistic visual experience

### v1

@@ 3.2 Cross-decomposition
The cross-covariance eigenspectrum for the testing set was then derived as:
@@
Statistical significance was assessed through block permutation tests to maintain the temporal structure in fMRI data. We permuted 6-second blocks of movie timeseries and recomputed the cross-decomposition procedure to generate a null distribution.
@@
3.3 Reliable individual differences
To assess whether there is reliable individual variability along the ranks of these high-dimensional spectra, we constructed individual differences matrices (IDMs) that capture the pairwise similarities of subjects  (?). For each logarithmic bin $\kappa$ of latent dimensions, each cell of an IDM was computed from a pair of subjects $i$ and $j$ as:
@@ 3.3 Reliable individual differences
After regressing out head motion confounds (intersubject framewise-displacement correlations and median framewise-displacement similarities) from  (?), we assessed reliability by computing Spearman correlations between mean IDMs derived from even and odd movies:
@@
To estimate confidence intervals, we implemented a bootstrap procedure with 1,000 resamples. For each resample, we randomly selected 90% of subjects with replacement and recomputed IDM correlations, excluding self-similarity values  (?). This generated a correlation distribution for each rank bin, from which we derived 95% confidence intervals. We computed IDM correlations both before and after partialing out ISC effects, with FDR correction applied when comparing differences between these conditions.
@@
3.4 Neural-behavioral similarity analysis
To examine relationships between neural individual differences and subjective movie interpretations, we analyzed the free recall data from  (?). Following the original study, we used transcribed verbal descriptions collected after each movie viewing. These descriptions were encoded into a 512-dimensional semantic space using Google’s Universal Sentence Encoder  (?). We constructed behavioral similarity matrices by computing pairwise cosine similarities between participants’ semantic embeddings.
We examined whether pairs of subjects with similar neural representational structure showed similar movie interpretations. First, we regressed out head motion confounds from all matrices. Then, we computed Spearman correlations between the behavioral similarity matrices and neural IDMs across two specific ranges of latent dimensions (1-10 and 11-100). This analysis was performed both with the original neural IDMs and with IDMs after controlling for voxelwise ISC matrices through partial correlation, allowing us to assess whether fine-grained neural patterns specifically contributed to behavioral similarities.
@@ 3.6 Code availability
Figure S1: Statistical validation of high-dimensional shared neural responses during movie viewing. Permutation tests confirm reliable stimulus-related variance extends to over hundred dimensions. This analysis establishes the statistical reliability of the high-dimensional shared neural representations analyzed in the main text. Grey lines indicate statistical thresholds (99.7%, 95.5%, and 68.3% percentiles) of null distributions generated by randomly block-shuffling movie-evoked fMRI responses across timepoints. The null distributions also display scale-dependent decay because the analysis reports covariances—the distribution is centered at zero at each rank but reflects the scale of variance at that dimensional bin. Dark violin plots show the distribution of observed cross-validated covariance values, which remain significantly above the 99.9% percentile of the null distribution up to approximately 150 dimensions in all regions. The separation between observed data and null distributions confirms that stimulus-related information extends across many latent dimensions, beyond what would be expected by chance. From left to right: early visual, ventral temporal, lateral temporal, and posterior parietal-cingulate regions.
@@
Figure S3: Region-specific patterns of individual differences emerge at each dimensional range. (a) Correlation matrices show relationships between individual differences matrices (IDMs) across regions and dimensional ranges. Each element represents the correlation between even-movie IDMs from one range in a region (rows) and odd-movie IDMs from another range in a region (columns). The log-scaled axis labels indicate dimensional ranges. The diagonal blocks show within-region correlations across matching dimensional ranges, while off-diagonal blocks show cross-region correlations. (b) Same analysis after controlling for coarse-scale effects by partialing out voxelwise inter-subject correlations. The more pronounced diagonal structure demonstrates that removing coarse-grained patterns reveals region-specific patterns of individual variability that are unique to each cortical area and dimensional range. This analysis complements Fig. 3 by showing that individual differences not only vary across dimensions within regions but also exhibit distinct patterns across regions at corresponding dimensional ranges.

### v2

@@ 3.4.1 Estimating latent dimension
This singular value decomposition performs the Procrustes transformation, identifying the optimal orthogonal rotations ($U$ for subject $i$, $V$ for subject $j$) that align the two subjects’ neural response patterns into a shared representational space that maximizes cross-covariance. The matrices $U$ and $V$ define the functional alignment between subjects. Then we projectected held-out test data onto the aligned dimensions (defined by $U$ and $V$ from training). The cross-covariance eigenspectrum for the testing set was then derived as:
@@
The test-set covariance spectrum, $\Sigma_{test}$, quantifies reliable stimulus-driven variance across dimension ranks. This test-set spectrum estimates dimensionality by identifying which dimensions carry reliable stimulus-driven signal as opposed to noise.
@@
Statistical significance was assessed through block permutation tests to maintain the temporal structure in fMRI data. We permuted 6-second blocks of test movie timeseries independently for both subjects and recomputed the cross-decomposition procedure to generate a null distribution from temporally desynchronized responses. This null distribution preserves the temporal autocorrelation structure of the data and thus provides a stringent test for whether observed covariance reflects meaningful stimulus-locked neural alignment. For each subject pair, we computed covariance spectra averaged across test movies and binned using the same logarithmic binning as observed data (n=1,000). At each dimensional bin, significance was assessed by comparing observed covariance against the permutation distribution ($p<0.05$, Bonferroni-corrected). For Fig. 1D, we show 10 individual permutation iteration traces from a randomly chosen single subject pair, along with the mean across all 1,000 permuted spectra for all subject pairs to illustrate the average null distribution centered at zero.
Random rotation testing
We performed additional random rotation control analyses to test whether any arbitrary rotation could produce the observed covariance pattern, as opposed to the specific Procrustes alignment that maximizes shared covariance. For each subject pair and test movie, we generated random orthogonal projections of the neural data and computed null covariance spectra. We generated random orthogonal projection matrices $Q_{x}\in\mathbb{R}^{v\times k}$ and $Q_{y}\in\mathbb{R}^{v\times k}$ for subjects $x$ and $y$, where $v$ is the number of voxels and $k$ is the number of components (set to maximum observed rank). These projection matrices were generated using QR decomposition of random Gaussian matrices. We projected each subject’s standardized data into these random $k$-dimensional subspaces ($X_{rot}=X_{test}Q_{x}$, $Y_{rot}=Y_{test}Q_{y}$) and computed covariance spectra from the randomly projected data. We found that the rotation null distributions converged quickly, and because these analyses were computationally intensive, we decided to stop the procedure after 100 iterations. On each iteration, the covariance statistics were movie-averaged, binned, and aggregated across subject pairs as with the observed data. Significance was assessed by comparing observed covariance against the random rotation distribution ($p<0.05$, Bonferroni-corrected).
@@
3.4.2 Reliable individual differences
Individual difference matrices
To assess whether there is reliable individual variability along the ranks of these high-dimensional spectra, we constructed individual differences matrices (IDMs) that capture the pairwise similarities of subjects [feilong2018reliable]. For each logarithmic bin $\kappa$ of latent dimensions, each cell of an IDM was computed from a pair of subjects $i$ and $j$ as:
@@ 3.4.2 Reliable individual differences
After regressing out head motion confounds (intersubject framewise-displacement correlations and median framewise-displacement similarities) from [savasegal], we assessed reliability by computing Spearman correlations between mean IDMs derived from even and odd movies:
@@
To ensure full independence between training and test data, we used a split-half approach where singular vectors were learned from even movies and applied separately to each odd movie to compute their IDMs. Reliability was assessed by correlating IDMs from different odd movies. We repeated this with train and test folds reversed and reported the average of the correlations across both folds. For the whole-brain analysis, we reported results from the even-fold test set only due to computational constraints.
@@
To estimate confidence intervals, we implemented a bootstrap procedure with 1,000 resamples. For each resample, we randomly selected 90% of subjects with replacement and recomputed IDM correlations, excluding self-similarity values [kriegeskorte2008representational]. This generated a correlation distribution for each rank bin, from which we derived 95% confidence intervals. We computed IDM correlations both before and after partialing out ISC effects, with FDR correction applied when comparing differences between these conditions.
Controlling for lower-rank dimensions
To assess whether higher-rank dimensions contain unique reliable information beyond what is captured by lower-rank dimensions, we performed successive partial correlation analyses. For each dimensional bin $c>1$, we sequentially regressed out IDMs from bins 1,,c11,...,c-1 from both the even- and odd-movie IDMs at bin $c$ in an iterative manner. We then computed the Spearman correlation between the final residual even- and odd-movie IDMs. Head motion confounds were regressed out prior to these analyses. Statistical significance was assessed using the same bootstrap procedure described above (1,000 resamples, 95% confidence intervals, Bonferroni correction).
@@
3.4.3 Neural-behavioral similarity analysis
To examine relationships between neural individual differences and subjective movie interpretations, we analyzed the free recall data from Sava-Segal et al. [savasegal] Following the original study, we used transcribed verbal descriptions collected after each movie viewing, and we encoded these descriptions into a 512-dimensional semantic space using Google’s Universal Sentence Encoder [cer2018universal]. We constructed behavioral similarity matrices by computing pairwise cosine similarities between participants’ semantic embeddings. Seva-Segal et al. [savasegal] selected USE because it was specifically trained to identify similarities between sentence pairs and because it best differentiated recall content between movies compared to alternative sentence encoders (BERT BASE, MiniLM, and MPNet). We repeated our key analyses using these alternative encoders and confirmed that the results were qualitatively similar.
We examined whether pairs of subjects with similar neural representational structure showed similar movie interpretations. First, we regressed out head motion confounds from all matrices. Then, we computed Spearman correlations between the behavioral similarity matrices and neural IDMs across different ranges of latent dimensions. This analysis was performed both with the original neural IDMs and with IDMs after controlling for voxelwise ISC matrices through partial correlation, allowing us to assess whether fine-grained neural patterns specifically contributed to behavioral similarities.
@@ 3.4.3 Neural-behavioral similarity analysis
Controlling for other dimensional ranges
To assess whether different dimensional ranges contain unique information about behavioral similarity beyond other ranges, we performed successive semi-partial correlation analyses. For each of the three dimensional ranges (1-10, 11-100, and 101-1000), we iteratively regressed out the neural IDMs from the other two dimensional ranges from the neural IDM of interest. Specifically, for a given dimensional range, we sequentially removed the influence of the other two ranges: we first regressed out the neural IDM from one of the other ranges to obtain a residual neural IDM, then regressed out the neural IDM from the remaining range from this residual. We then computed the Spearman correlation between the final residual neural IDM and the behavioral IDM. Head motion confounds were regressed out from all neural IDMs prior to these analyses. We also performed this analysis after controlling for voxelwise ISC matrices (as described above). The same permutation testing procedure (1,000 iterations, FDR correction) was applied to assess statistical significance.
3.4.4 Concreteness and shared neural covariance
Mean covariance analysis
To relate the recall narrative concreteness to neural patterns across dimensional bins (ranks 1-10, 11-100, 101-1000), we computed each subject’s mean neural covariance for each dimensional bin by averaging all IDM values involving that subject (41 pairwise comparisons per subject given n=42 subjects). We then computed the Spearman correlations between these subject-level neural values and the recall narrative concreteness scores (See ”Narrative concreteness”). Statistical significance was assessed through permutation tests with 1,000 iterations, FDR-corrected. To isolate unique variance per bin, we additionally performed successive semi-partial correlations controlling for the other two bins, iteratively removing their effects before computing correlations with concreteness scores (Fig. S5).
Dimensionality count analysis
We quantified individual neural dimensionality by counting, for each subject, how many dimensions showed significant covariance (permutation test, 99.5th percentile threshold) across all subject pairs involving that subject, evaluating the significance of each dimension without binning. We then computed the subject-level dimensionality counts averaged across movies and computed the Spearman correlation with the recall concreteness scores.
@@ 3.6 Code availability
Figure S2: High-dimensional individual differences remain reliable after controlling for low-dimensional structure. Split-half reliability of IDMs observed across dimensional ranges in Fig. 2B persists even after sequentially removing contributions from lower-dimensional ranges. Dark shades represent correlations between original IDMs; lighter shades show correlations after controlling for coarse-scale effects through partialing out voxelwise ISC IDMs. Grey horizontal line indicates zero correlation. Dots indicate statistical significance ($p<0.05$, Bonferroni-corrected); bootstrap confidence intervals from 1,000 resamples. From left to right: early visual, ventral temporal, lateral temporal, and posterior parietal-cingulate regions.
Figure S3: Region-specific patterns of individual differences emerge at each dimensional range. (a) Correlation matrices show relationships between individual differences matrices (IDMs) across regions and dimensional ranges. Each element represents the correlation between even-movie IDMs from one range in a region (rows) and odd-movie IDMs from another range in a region (columns), normalized by the corresponding diagonal reliabilities such that diagonal values equal 1. The log-scaled axis labels indicate dimensional ranges. The diagonal blocks show within-region correlations across matching dimensional ranges, while off-diagonal blocks show cross-region correlations. (b) Same analysis after controlling for coarse-scale effects by partialing out voxelwise inter-subject correlations. The more pronounced diagonal structure demonstrates that removing coarse-grained patterns reveals region-specific patterns of individual variability that are unique to each cortical area and dimensional range. This analysis complements Fig. 3 by showing that individual differences not only vary across dimensions within regions but also exhibit distinct patterns across regions at corresponding dimensional ranges.
Figure S4: High-dimensional individual differences are strongest in sensory and semantic cortices. (A) Maps show correlation across fine-grained brain parcellations for dimensions 1-10 (left), 11-100 (middle), and 101-1000 (right). Individual differences extending to high-dimensional ranges are strongest in sensory regions (visual and auditory cortices) and semantic/default mode regions (precuneus, angular gyrus), modest in memory-related regions (hippocampus, medial prefrontal cortex), and diminished in other regions. This spatial distribution indicates that multidimensional individual differences reflect meaningful stimulus-driven variability concentrated in regions engaged by naturalistic audiovisual content. Maps show raw correlations between odd-movie IDMs without thresholding. Black outlines indicate main analysis ROIs.
Figure S5: Unique contributions of different dimensional ranges to behavioral predictions. (A) Partial correlations between neural IDMs and behavioral similarity, controlling for all other dimensional ranges, to supplement Fig. 4B. (B) Partial correlations between shared neural variance and narrative concreteness, controlling for all other dimensional ranges, to supplement Fig. 5B. Left shows original neural covariance; Right shows results after removing coarse-scale effects. Bars show mean partial correlations; dots indicate statistical significance ($p<0.05$, FDR-corrected). Small symbols represent individual movies. While attenuated compared to zero-order correlations, several dimensional ranges maintained significant unique contributions.